OpenStudy (anonymous):

dentify the pattern for the following sequence. Find the next three terms in the sequence. -15, -11, -7, -3, ____, ____, ____,...

OpenStudy (gorv):



OpenStudy (adilalvi):

1,5,9,13,17,21 and so on adding +4.

OpenStudy (gorv):

in each term a fixed number is added ....can u guess the next 3 term by addin that fixed number to previous term

OpenStudy (anonymous):

1,5,9

OpenStudy (anonymous):

So you know it's a sequences right, So first of all find the d or common difference which is 4 because you add 4. So. -3 + 4 = 1 1 + 4 = 5 5+4 = 9

OpenStudy (anonymous):

By the way you could have used \[U _{n+1}- U _{n}=d\] This is just plugging in the first term: -15 for and the sequential terms